# JWST Exoplanetary Worlds and Elemental Survey (JEWELS) I: High-Precision Chemical Abundances of 20 FGK Planet-Hosting Stars from JWST Cycle 2

### v2

@@ Abstract
We present high-precision chemical abundances for 20 FGK stars hosting planets observed in JWST Cycle 2 GO programs. Using high-resolution, high–signal-to-noise ratio spectra from the ESO and Keck archives, we perform a strict line-by-line differential analysis relative to the Sun to derive stellar parameters and abundances of 19 elements from C to Zn. The stars span effective temperatures of 4500–6500 K and metallicities from –0.57 to +0.50 dex. The sample includes hosts of both gas giants and terrestrial planets, allowing direct comparison between stellar composition and planetary properties. Several of the giant planets orbit metal-rich stars. The detailed abundance patterns show clear chemical diversity, including carbon-enhanced but mildly metal-poor stars (TOI-824, TOI-1130, GJ 9827) and $\alpha$-enhanced metal-poor stars (TOI-561, GJ 9827, TOI-824). These variations trace differences in protoplanetary disk composition and may influence planetary interiors and atmospheric chemistry. The planet-hosts show a range of [C/O] ratios, and the diverse [Mg/Si] ratios may suggest varied interior compositions for their rocky planets. This homogeneous stellar abundance, together with future uniform JWST planetary atmosphere measurements, provides a foundation for exploring the planet mass–metallicity relation and the connection between stellar chemistry and planetary formation pathways. These results constitute the first step in a larger survey spanning multiple JWST cycles to systematically examine how host star composition shapes exoplanetary systems.
1 Introduction
@@
2 Target Selection and Observations
@@
3 Stellar Atmosphere Parameters
@@
4 Chemical Abundance Analysis
Elemental abundances were derived following the procedures described in Sun et al. (2025b, a). We adopted a strictly line-by-line differential abundance analysis relative to the solar spectrum to compute [X/H]222[X/H] = A(X) – A(X), where X denotes the element. A(X) = 12 + $\log(N_{\rm X}/N_{\rm H})$, and $N_{\rm X}$ and $N_{\rm H}$ are the number densities of element X and hydrogen, respectively.. The measured EWs were analyzed with the abfind driver in MOOG to derive individual line abundances, A(X), for each element. The selected lines include those important for tracing planetary atmospheric compositions. For oxygen, we applied the NLTE corrections from Ramírez et al. (2007); for all other elements, NLTE effects are expected to be negligible and largely cancel in the differential analysis.
@@ 4 Chemical Abundance Analysis
5 Discussion
@@
Beyond overall metallicity, detailed abundance patterns reveal additional diversity. The sample includes stars with near-solar compositions and others showing distinct chemical peculiarities, such as carbon-enhanced but slightly metal-poor stars (TOI-824, TOI-1130, GJ 9827), volatile-rich stars (WASP-69, WASP-52), and old, $\alpha$-enhanced, metal-poor stars (TOI-561, GJ 9827). These abundance differences trace the composition of the natal protoplanetary disks and can influence planet formation outcomes. For example, the stellar [C/O] affects the atmospheric chemistry of gas giants, while refractory element ratios such as [Mg/Si] shape the interior composition of terrestrial planets. Together, these stellar chemical signatures provide essential context for interpreting the structure, bulk properties, and atmospheres of the associated planetary systems.
@@
Figure 2: Relation between stellar chemistry and planetary systems within the JWST Cycle 2 GO sample. (a) $\alpha$-element abundance ratio versus metallicity, tracing Galactic chemical evolution. (b) Planetary radius versus stellar [C/O], where high [C/O] may indicate carbon-rich atmospheres. Both planets in the TOI-1130 system are shown and connected by a black dashed line. TOI-125 lacks an [O/H] measurement and is not shown. (c) Bulk density of rocky planets versus [Mg/Si]. The planets TOI-1130 b and TOI-125 b are included in this panel. (d) Stellar carbon enhancement ([C/Fe]) versus [Fe/H], highlighting carbon-rich metal-poor stars.
@@
Several stars in the sample show carbon enrichment (Figure 2d). GJ 9827, TOI-824, and TOI-1130 have [C/H] values of 1.011, 1.214, and 1.639 dex, respectively, despite having slightly subsolar metallicities ([Fe/H] = 0.612-0.612, 0.196-0.196, and 0.141-0.141 dex). This results in [C/Fe] >+1.4>+1.4 dex. One possible explanation is mass transfer from an evolved asymptotic giant branch (AGB) companion (e.g., Lucatello et al. 2005; Masseron et al. 2010). Enrichment from early supernovae is unlikely because these stars do not reach the very low metallicities of [Fe/H] <1.0<-1.0 dex. Apart from carbon-enriched metal-poor stars, WASP-69 also exhibits elevated C, O, and S abundances, indicative of a volatile-rich formation environment. We note, however, that the carbon abundances for TOI-1130, TOI-824, GJ 9827, and WASP-69 may be less reliable because only a few carbon lines are measurable in cooler dwarfs with $T_{\rm eff}<5000$ K. Consequently, the inferred abundances likely carry substantial uncertainties. We also examined several carbon-bearing molecular features and synthesized both individual lines and molecular bands, but the synthetic spectra reproduce neither the carbon features nor the surrounding wavelength regions well. This includes the C2 bands at 5128 Å and 5165 Å, as discussed in da Silva et al. (2024).
@@
To examine whether mass transfer is feasible for GJ 9827, TOI-824, and TOI-1130, we first check for evidence of companions. As JWST targets, these stars have extensive RV follow-up, which should reveal any nearby massive companion. The literature shows no sign of a secondary object around TOI-824 or GJ-9827. For TOI-1130, Korth et al. (2023b) report a linear RV trend that may indicate an outer massive companion, though this has not yet been confirmed. We next examine the abundances of s-process elements (Sr, Y, Ba), which can indicate past AGB mass transfer. We perform spectrum synthesis using the synth driver in MOOG for Sr, Y, and Ba, and apply the same procedure to a solar spectrum taken with the same instrument. All three stars show strong Sr absorption lines, with EWs more than twice those of the Sun, indicating significant Sr enrichment. The Y and Ba lines are also detected though in general weaker than the solar lines. Table 3 lists the synthesized abundances for each line and the mean abundances calculated in linear space. The uncertainties are estimated in the same way as in previous sections, coadding the propagated atmospheric-parameter errors and $\sigma_{\mu}$ in quadrature. In all three stars, Sr is above solar and strongly enhanced relative to iron, with [Sr/Fe] values of 1.49, 1.07, and 0.91 dex. In contrast, Y and Ba are below solar in all cases, but their absolute abundances are consistent with the stars’ slightly subsolar metallicities. Although Sr is enhanced, the lack of corresponding Y and Ba enhancement may not suggest mass transfer from an AGB companion.
@@
In Figure 2b, planet radius is plotted against [C/O]. The [C/O] values vary widely across systems; in particular, the multi-planet system TOI-1130 exhibits very high [C/O]. The [Mg/Si] ratio also provides insight into the composition of rocky planets. As shown in Figure 2c, stars in the sample span a range of [Mg/Si]. This ratio controls the dominant mineralogy of terrestrial mantles (Bond et al., 2010; Delgado Mena et al., 2010): low [Mg/Si] values favor feldspar-rich silicates, while high [Mg/Si] values produce olivine- and pyroxene-dominated interiors. Variations in this ratio suggest that rocky planets around different hosts may have diverse interior structures, mantle convection patterns, and volcanic activity, resulting in a wider range of geological and thermal behaviors than found in the Solar System.
@@ acknowledgements
This work is supported by the Science and Technology Commission of Shanghai Municipality under Grant No. 25ZR1402244, the National Key R&D Program of China under Grant No. 2024YFA1611801, and the Shanghai Jiao Tong University Funds Program No. AF4260012.

### v3

@@ Abstract
We present high-precision chemical abundances for 20 FGK stars hosting planets observed in JWST Cycle 2 GO programs. Using high-resolution, high–signal-to-noise ratio spectra from the ESO and Keck archives, we perform a strict line-by-line differential analysis relative to the Sun to derive stellar parameters and abundances of 19 elements from C to Zn. The stars span effective temperatures of 4500–6500 K and metallicities from –0.57 to +0.50 dex. The sample includes hosts of both gas giants and terrestrial planets, allowing direct comparison between stellar composition and planetary properties. Several of the giant planets orbit metal-rich stars. The detailed abundance patterns show clear chemical diversity, including carbon-enhanced but mildly metal-poor stars (TOI-824, TOI-561, TOI-1130, GJ 9827) and $\alpha$-enhanced metal-poor stars (TOI-561, GJ 9827, TOI-824). These variations trace differences in protoplanetary disk composition and may influence planetary interiors and atmospheric chemistry. The planet-hosts show a range of [C/O] ratios, and the diverse [Mg/Si] ratios may suggest varied interior compositions for their rocky planets. This homogeneous stellar abundance, together with future uniform JWST planetary atmosphere measurements, provides a foundation for exploring the planet mass–metallicity relation and the connection between stellar chemistry and planetary formation pathways. These results constitute the first step in a larger survey spanning multiple JWST cycles to systematically examine how host star composition shapes exoplanetary systems.
I Introduction
@@
II Target Selection and Observations
@@
III Stellar Atmosphere Parameters
@@
IV Chemical Abundance Analysis
Elemental abundances were derived following the procedures described in Sun et al. (2025b, a). We adopted a strictly line-by-line differential abundance analysis relative to the solar spectrum to compute [X/H]222[X/H] = A(X) – A(X), where X denotes the element. A(X) = 12 + $\log(N_{\rm X}/N_{\rm H})$, and $N_{\rm X}$ and $N_{\rm H}$ are the number densities of element X and hydrogen, respectively.. The measured EWs were analyzed with the abfind driver in MOOG to derive individual line abundances, A(X), for each element. The selected lines include those important for tracing planetary atmospheric compositions. For oxygen, we applied the NLTE corrections from Ramírez et al. (2007); for all other elements, NLTE effects are expected to be small and largely cancel in the differential analysis.
@@ IV Chemical Abundance Analysis
As noted in previous work (e.g. da Silva et al. 2024), for dwarfs cooler than 5000 K, atomic carbon lines are often too weak to be measured confidently or are strongly blended with molecular features. We observe the same behavior in our sample: for the four cool dwarfs TOI-1130, TOI-824, GJ 9827, and WASP-69, carbon abundances derived from atomic-line EWs are too high. We therefore derive carbon abundances by synthesizing the C2 molecular bands at 5165 Å (da Silva et al. 2024) using the synth driver in MOOG. The adopted abundances in Table 2 correspond to the best-fitting synthetic spectra, which simultaneously reproduce the C2 band features and the surrounding continuum and neighboring lines.
V Discussion
@@
Beyond overall metallicity, detailed abundance patterns reveal additional diversity. The sample includes stars with near-solar compositions and others showing distinct chemical peculiarities, such as carbon-enhanced but slightly metal-poor stars (TOI-824, TOI-561, TOI-1130, GJ 9827), volatile-rich stars (WASP-69, WASP-52), and old, $\alpha$-enhanced, metal-poor stars (TOI-561, GJ 9827). The overall $\alpha$-element abundance is computed as the mean of the Mg, Si, Ca, and Ti abundances, where measurements for these elements are available. These abundance differences trace the composition of the natal protoplanetary disks and can influence planet formation outcomes. For example, the stellar [C/O] affects the atmospheric chemistry of gas giants, while refractory element ratios such as [Mg/Si] shape the interior composition of terrestrial planets. Together, these stellar chemical signatures provide essential context for interpreting the structure, bulk properties, and atmospheres of the associated planetary systems.
@@
Figure 2: Relation between stellar chemistry and planetary systems within the JWST Cycle 2 GO sample. (a) $\alpha$-element abundance ratio versus metallicity, tracing Galactic chemical evolution. (b) Planetary radius versus stellar [C/O], where high [C/O] may indicate carbon-rich atmospheres. Both planets in the TOI-1130 system are shown and connected by a black dashed line. TOI-125 lacks an [O/H] measurement and is not shown. (c) Bulk density of rocky planets versus [Mg/Si]. The planets TOI-1130 b and TOI-125 b are included in this panel. (d) Stellar carbon enhancement ([C/Fe]) versus [Fe/H], highlighting carbon-rich while slightly metal-poor stars.
@@
Several stars in the sample show carbon enrichment (Figure 2d). GJ 9827, TOI-561, TOI-824, and TOI-1130 have [C/Fe] values of 0.23, 0.291, 0.216, and 0.241 dex, respectively, despite having slightly subsolar metallicities ([Fe/H] = -0.612, -0.379, -0.196, and -0.141 dex). One possible explanation is mass transfer from an evolved asymptotic giant branch (AGB) companion (e.g., Lucatello et al. 2005; Masseron et al. 2010). Enrichment from early supernovae is unlikely because these stars do not reach the very low metallicities of [Fe/H] <1.0<-1.0 dex. Apart from carbon-enriched metal-poor stars, WASP-69 also exhibits elevated C, O, and S abundances, indicative of a volatile-rich formation environment.
@@
To examine whether mass transfer is feasible for GJ 9827, TOI-561, TOI-824, and TOI-1130, we first check for evidence of companions. As JWST targets, these stars have extensive RV follow-up, which should reveal any nearby massive companion. The literature shows no sign of a secondary object around TOI-824, TOI-561, or GJ-9827. For TOI-1130, Korth et al. (2023b) report a linear RV trend that may indicate an outer massive companion, though this has not yet been confirmed. We next examine the abundances of s-process elements (Sr, Y, Ba), which can indicate past AGB mass transfer. We perform spectrum synthesis using the synth driver in MOOG for Sr, Y, and Ba, and apply the same procedure to a solar spectrum taken with the same instrument. Apart from TOI-561, the rest three stars show strong Sr absorption lines, with EWs more than twice those of the Sun, indicating significant Sr enrichment. The Y and Ba lines are also detected though in general weaker than the solar lines. Table 3 lists the synthesized abundances for each line and the mean abundances calculated in linear space for GJ 9827, TOI-824, and TOI-1130. The uncertainties are estimated in the same way as in previous sections, coadding the propagated atmospheric-parameter errors and $\sigma_{\mu}$ in quadrature. For GJ 9827, TOI-824, and TOI-1130, Sr is above solar and strongly enhanced relative to iron, with [Sr/Fe] values of 1.49, 1.07, and 0.91 dex. In contrast, Y and Ba are below solar in all cases, but their absolute abundances are consistent with the stars’ slightly subsolar metallicities. Although Sr is enhanced, the lack of corresponding Y and Ba enhancement may not suggest mass transfer from an AGB companion.
@@
In Figure 2b, planet radius is plotted against [C/O]. The [C/O] values vary across systems; in particular, the multi-planet system TOI-1130 exhibits high [C/O]. The [Mg/Si] ratio also provides insight into the composition of rocky planets. As shown in Figure 2c, stars in the sample span a range of [Mg/Si]. This ratio controls the dominant mineralogy of terrestrial mantles (Bond et al., 2010; Delgado Mena et al., 2010): low [Mg/Si] values favor feldspar-rich silicates, while high [Mg/Si] values produce olivine- and pyroxene-dominated interiors. Variations in this ratio suggest that rocky planets around different hosts may have diverse interior structures, mantle convection patterns, and volcanic activity, resulting in a wider range of geological and thermal behaviors than found in the Solar System.
@@ acknowledgements
We thank Elisa Delgado Mena for helpful discussions regarding carbon abundances. This work is supported by the National Key R&D Program of China under Grant No. 2024YFA1611801, the Science and Technology Commission of Shanghai Municipality under Grant No. 25ZR1402244, and the Shanghai Jiao Tong University Funds Program No. AF4260012. This work is also supported in part by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 24DX1400100, ZJ2023-ZD-001). This work is based, in part, on observations made with the NASA/ESA/CSA James Webb Space Telescope.